OpenStudy (anonymous):

find the first term of the arithmetic sequence having the data given. an = 48, Sn = 546, n = 26

OpenStudy (saifoo.khan):

Do you know the formula?

OpenStudy (anonymous):

ish... not that well. Im just starting this and I just need some examples done.

OpenStudy (anonymous):

\[\S_{n}=\frac{n}{2}(a_{1}+a_{n})\]

OpenStudy (anonymous):

Sum.. \[\LARGE Sn=\frac n2 [2a_1+(n-1)\cdot d]\] \[\LARGE a_n=a_1+(n-1)\cdot d\]